OpenStudy (anonymous):

Solve. 4x^2 + 8x = 252 x = 7 and x = 9 x = -7 and x = 9 x = -7 and x = -9 x = 7 and x = -9

OpenStudy (anonymous):

x = 7 and x = -9

OpenStudy (anonymous):

x^2+2x=63 x^2+2x+1=63+1 (x+1)^2=8^2 x+1=+/-8 x=7/x=-9